### v1

# A Fast Relax-and-Round Approach to Unit Commitment with Sub Hourly Mechanical and Ramp Constraints

Abstract
We propose a novel computational method for unit commitment (UC), which does not require linearized approximation and provides several orders of magnitude performance improvement over current state-of-the-art. The performance improvement is achieved by introducing a heuristic tailored for UC problems. The method can be implemented using existing continuous optimization solvers and adapted for different applications. We demonstrate value of the new method in examples of advanced UC analyses at the scale where use of current state-of-the-art tools is infeasible. We expect that the capability demonstrated in this paper will be critical to address emerging power systems challenges with more volatile large loads, such as data centers, and generation that is composed of larger number of smaller units, including significant behind-the-meter generation.
I Introduction
Recent changes in the electric power industry indicate a growth in load, a decrease in the average size of generating units, and an increasing complexity in forecasting. According to the U.S. Energy Information Administration (EIA), over the past 50 years total electricity production across all sectors increased by a factor of 2.3, from 1.87 thousand TWh in 1974 to 4.3 thousand TWh in 2024 [34].
This growth is expected to intensify due to the increasing digital share of the US economy. For instance, 2024 S&P Global Market Intelligence Datacenters and Energy Report indicates that 43.8 GW of datacenter capacity are planned for deployment between 2025 and 2030[29]. At nominal capacity, data centers could contribute up to 384 TWh of annual electricity consumption, almost a 10% increase in load from a single load category.
Generator EIA data shows they are becoming smaller. For units in operation in 2024, excluding Alaska and Hawaii, average nameplate capacity for a generator built in 1974 was 205.22 MW. For a generator built in 2024 it was 49.5 MW [8]. For 2408 generating units listed in “proposed”, the average nameplate capacity was 80.21 MW for units dated 2014-2024 and 193.83 MW for units planned for 2025-2030. However, the actual average capacity is expected to be lower. First, historically, proposed units between 2014 and 2024 were approximately 1.5-4.5 times larger than those ultimately built.
Furthermore, EIA statistics can be assumed to be biased towards larger generating units belonging to utilities. A large part of new generating units may be coming from behind-the-meter generation and may be significantly smaller [11]. A basic assessment of the reviewed data shows that if the electricity demand grew by a factor of nearly 2.5 and the average generating unit size decreased by a factor of 4, the total number of generating units in the system, including those not accounted for by EIA, could increase by a factor of 10, based on past trends, and this trend is likely to continue.
In addition to size, there remains a question of time related system complexity. According to EIA, the share of generation from all thermal sources, which are traditionally considered predictable, increased from 83.7% in 1974 to 93.1% in 2004, but declined to 78.9% in 2024 [34]. This decline, were it to continue, would make the generation stack less predictable, contributing to increased operational uncertainty and greater computational and modeling complexity.
Intra-hour constraints are becoming significantly more important, as evidenced by the IEEE Task Force on Solving Large Scale Optimization Problems in Electricity Market and Power System Applications [6]. However, higher granularity of unit commitment involves more frequent unit commitment calculation: 4 times more frequent for the 15 minute interval currently used by many researchers, and 12 time more frequent for the 5 minute interval, which is often used for modeling non thermal resources.
Meanwhile, unit commitment formulations and solvers remain limited. Recent literature reviews show system sizes for existing academic studies, which tend to use CPLEX or Gurobi on standard production-grade hardware. The order of magnitude of the problem is usually in $100$s of generators (not buses). For instance, [6] discusses one study with 158 units, [10] examines studies with 20-130 units, [19] surveys studies with 5-1870 units, with most of references in the range of 10-200 units, [14] reviews systems with a stated system size of up to 100 units. A survey of large unit commitment problems of 5-2709 generators is discussed in [35], with most reviewed studies using 100-200 units (Note: the original paper referenced by [35] uses 2704 generators [12]). The studies directly targeting systems with finer temporal granularity show very limited scale: [23] discusses a system with 32 generators for 30-minute unit commitment, [21] uses 54 units, 15-minute steps, [15] uses varying extents of granularity up to 5-minute intervals on a system with 10 dispatchable generating units.
Studies directed at improving computational efficiency use comparably small simulations and one hour resolution. For example, [16] uses a bundle method on a 130-generator system, [37] uses temporal bundles and 289 generators, and [30] demonstrates security constrained unit commitment on 69 generators. There are several reasons for the use of small system sizes in unit commitment problems. First, system scaling happens along both spatial and temporal granularity. Running 1000 unit commitment scenarios for a 150-unit system remains substantially faster than running one scenario for a 1000-unit system. Similarly, running 72 time decoupled unit commitment scenarios is faster than running one time coupled scenario for a three days outlook. Second, existing unit commitment algorithms rely on mixed integer programs (MIPs) [2]. In the worst case, solving MIPs is NP-hard and requires time exponential in the problem size, to enumerate all possibilities. For more information about the mathematical reasons for the poor scaling behavior of MIPs, see optimization-specific reviews such as [17, 20, 4, 28].
In earlier work [27], we proposed a fast approach to unit commitment for small, fast-ramping generating units, which are expected to play an increasingly important role due to the rapid expansion of data center infrastructure. In this study, we extend this approach to include larger and slower generating units found in today’s conventional grids. To achieve this, we first review the basic UC formulation from [27] and then augment it with supplementary constraints usually found in literature [6, 19].
Our approach only accounts for the costs of generation in the current period, but includes intertemporal runtime constraints and, later, ramping constraints. From that perspective, the algorithm is partially temporally decoupled and cannot be directly compared to fully coupled UC algorithms. However, it still presents an advancement over earlier methods because (i) it scales sub-quadratically rather than exponentially and (ii) does not linearize the problem, making solutions more accurate. Each extended formulation is evaluated from two viewpoints: algorithmic runtime and the objective function. For the relaxed unit commitment formulation, we further analyze the stochastic characteristics of the objective function.
The study is organized as follows. Section II summarizes the original model. Section III introduces intraday unit constraints. Section IV extends the model to include ramping constraints. Section V briefly discusses results with a different ramping formulation. Section VI summarizes the results and suggests further research directions.
II Review of Original Unit Commitment Problem
II-A Mathematical Formulation
In our previous work [27], the problem formulation was designed to address a very specific system need, i.e. to accommodate a potentially large number of small- to medium-sized flexible resources, potentially with highly convex supply curves. This resulted in the objective function that emphasized cost structure compared to equipment or grid constraints. Each unit $i$ was modeled with quadratic cost curves and allowed to commit up to a certain volume between $P_{\min,i}$ and $P_{\max,i}$ as opposed to committing the entire $P_{\max,i}$ of a unit.
To address possible demand uncertainties, we introduce the reserve margin, similar to the robust approach discussed in literature [19, 21, 38]. This ensures that if our largest committed generator goes out, and demand is higher than predicted, we can still supply it. Here we choose a simple Gaussian noise model, and will build on this in future work.
The respective optimization problem for $n$ generators is:
$\min\quad$ $\sum_{i=1}^{n}\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)u_{i}$ (1a)
s.t. $P_{\min,i}\leq P_{i}\leq P_{\max,i},\;u_{i}\in\{0,1\}$ (1b)
$\sum_{i=1}^{n}u_{i}P_{i}\geq D,$ (1c)
$\sum_{i=1}^{n}u_{i}P_{\max,i}\geq D+3\sigma_{D}+R$ (1d)
The parameters supplied to the optimization solver are:
  • $a_{i},b_{i},c_{i}$: quadratic cost coefficients for each generator $i$.
  • $P_{\max,i},P_{\min,i}$: maximum and minimum power output for generator $i$.
  • $D$: current predicted demand.
  • $\sigma_{D}$: standard deviation of demand forecast
  • $R=\max\{P_{\max,i}\}$: reserve margin.
The decision variables are:
  • $P_{i}\in[P_{\min,i},P_{\max,i}]$ for $1\leq i\leq n$: power output of each generator $i$.
  • $u_{i}\in\{0,1\}$ for $1\leq i\leq n$: commitment status of generator $i$.
The constraint Eq. 1c ensures that we supply the projected demand. The constraint Eq. 1d ensures that if our largest generator goes offline we will be able to supply demand if it is within 3 standard deviations of our prediction. This ensures we can still supply demand if any one generator goes out and our predicted demand is higher than expected. More complicated uncertainty models can fit into this framework, but developing them is beyond the scope of this work.
We solve this problem by first relaxing $u_{i}\in\{0,1\}$ to $y_{i}\in[0,1]$ and solving the relaxed continuous problem. We reorder terms in the descending order in $y_{j}$ so that $y_{j}>y_{j+1}$. Then, we find the first $m$ generators such that $\sum_{j=1}^{m}P_{\max,j}\geq D+3\sigma_{D}+\max_{j}[P_{\max,j}]$, and solve economic dispatch with all $m\leq k\leq n$, for a total of nm+1n-m+1. The solution is the subproblem that produced the lowest objective.
II-B Simulation and Results
Figure 1: Test results for 46-1840 generating units [27]. (Top) convergence time for both solvers, RRUC outperforms MIP by orders of magnitude. (Bottom) deviation in objective value calculated as (RRUC-MIP)/MIP. For larger problems, the deviation is below the tolerance set for the solvers.
To verify the performance of the approach, we used an increasingly large single node system with varied generators obtained from PJM Data Miner [24]. We fitted the piece-wise linear bids from PJM to quadratic cost curves with least squares, in line with [36, 1].
We implemented our method in Julia [5] with the MadNLP optimization package [22], which is based on our prior work [26, 32]. As a reference MIP implementation, we choose the open-source Juniper package [18], which runs on the same JuMP platform [7] as MadNLP. We ran both solvers on an Apple M3 Pro Laptop with 18GB RAM.
We highlight the key results of [27] in Fig. 1. Relax-and-Round Unit Commitment (RRUC) was $1,000$s of times faster than the MIP solver reference. We abandoned the latter after 270 units as it was taking about 10 hours on the available computational resources.
The accuracy of the RRUC algorithm compared to MIP depended on the system size. On very small problems, a traditional MIP solver committed different units and achieved a smaller cost (objective function). The difference was under $6\%$. However, the difference decreased with the increasing problem size. On the 270-generator problem, the largest problem the MIP solver could solve, the difference in objective function was smaller than the tolerance set by the solvers and the two solvers selected the same units.
III Unit Commitment with Runtime Constraints
III-A Mathematical Formulation
We now expand the original formulation to include a wider variety of generators, some of which are slower large units found in utility power plants. Such units have runtime constraints and a maximum number of daily starts. Their startup and shutdown costs are a significant portion of their total operating costs. The different components of runtime and other mechanical constraints are summarized in  [35, 36, 33].
We follow the general consensus using startup and shutdown costs and minimum on/off time constraints. Specifically, for each generator we add a minimum runtime $T_{\min,i}$ and a preprocessing stage before optimization. The preprocessing determines how long a generator has been running. If a generator is on but has run fewer periods than $T_{\min,i}$, the generator cannot be switched off. Such generators are considered must run and denoted as group $\mathcal{G}_{m}$.
Further, each generator has a maximum number of startups during the 24-hour window, $S_{\max,i}$. The generators that are off and cannot yet be switched on due to the maximum number of starts are excluded from the optimization problem for that time period. The remaining generators are discretionary. We denote this group $\mathcal{G}_{d}$. The resulting optimization problem is:
$\min\quad$ $\sum_{i\in G_{m}}\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)$
$+\sum_{i\in G_{d}}\left[\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)u_{i}+K_{i}(u_{i}-u_{i}^{0})^{2}\right]$ (2a)
s.t. $P_{\min,i}\leq P_{i}\leq P_{\max,i},\;u_{i}\in\{0,1\}$ (2b)
$\sum_{i\in G_{m}}P_{i}+\sum_{i\in G_{d}}u_{i}P_{i}\geq D,$ (2c)
$\sum_{i\in G_{m}}P_{\max,i}+\sum_{i\in G_{d}}u_{i}P_{\max,i}\geq{D}_{\max,72}+3\sigma_{D}+R,$ (2d)
$\sum_{i\in G_{m}}P_{\min,i}+\sum_{i\in G_{d}}u_{i}P_{\min,i}\leq{D}_{\min,72}-3\sigma_{D}.$ (2e)
The parameters supplied to the optimization solver are:
  • $a_{i},b_{i},c_{i}$: quadratic cost coefficients usually resulting from fuel costs or strategic bidding behavior.
  • $P_{\max,i},P_{\min,i}$: maximum capacity and minimum stable output, usually determined by mechanical characteristics of the units.
  • $D$, $\sigma_{D}$: demand volume and the possible demand deviations in the robust model formulation.
  • $K_{i}$: commitment change penalty for generator $i$. The true costs of startup and shutdown are usually not the same. Averaging the startup and shutdown costs prevents market distortions and ensures that there are no artificial incentives to run.
  • $u_{i}^{0}\in\{0,1\}$: previous commitment status of generator $i$
  • ${D}_{\min,72}$ and ${D}_{\max,72}$: minimum and maximum volumes of predicted demand in next 72 hours.
  • $\mathcal{G}_{m}$: the set of must run generators.
  • $\mathcal{G}_{d}$: the set of discretionary generators.
  • $R=\max\{\max_{i\in\mathcal{G}_{m}}P_{\max,i},\max_{i\in\mathcal{G}_{d}}P_{\max,i}\}$: reserve margin (changed to only include generators that can run).
The decision variables are:
  • $P_{i}\in[P_{\min,i},P_{\max,i}]$ for $i\in\mathcal{G}_{m}$: power output of must-run generator $i$.
  • $P_{i}\in[P_{\min,i},P_{\max,i}]$ for $i\in\mathcal{G}_{d}$: power output of discretionary generator $i$.
  • $u_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$: commitment status of discretionary generator $i$.
Similar to the problem Section II, the problem defined in Eq. 2 is a MIP, an NP-hard type problem. This problem is intractable and usually solved by a branch and bound method [20]. We relax $u_{i}\in\{0,1\}$ to $y_{i}\in[0,1]$, reorder terms in the descending order, and determine which units are preferred to run. Since in our problem the $y_{i}$ variables represent which generators are committed and which are not, we find the first $\tilde{m}$ generators such that
$\sum_{i\in\mathcal{G}_{m}}P_{\max,i}+\sum_{\begin{subarray}{c}1\leqslant i\leqslant\tilde{m}\\ i\in\mathcal{G}_{d}\end{subarray}}P_{\max,i}\;\geq\;D_{\max,72}+3\sigma_{D}+R_{\tilde{m}}$, (3)
where
$R_{\tilde{m}}=\max\!\left\{\max_{i\in\mathcal{G}_{m}}P_{\max,i},\;\max_{\begin{subarray}{c}1\leqslant i\leqslant\tilde{m}\\ i\in\mathcal{G}_{d}\end{subarray}}P_{\max,i}\right\}.$
As discussed above, some previously committed units must operate due to runtime constraints. Therefore, $m\doteq|\mathcal{G}_{m}|+\tilde{m}$. We can calculate $m_{\max}$, the maximum number of generators that can operate, using Eq. 2e analogously to the way we used Eq. 2d to derive Eq. 3. This gives us Eq. 4:
$\sum_{i\in\mathcal{G}_{m}}P_{\min,i}+\sum_{\begin{subarray}{c}1\leqslant i\leqslant\tilde{m}_{\max}\\ i\in\mathcal{G}_{d}\end{subarray}}P_{\min,i}\;\geq\;D_{\min,72}+3\sigma_{D},$ (4)
with $m_{\max}\doteq|\mathcal{G}_{m}|+\tilde{m}_{\max}$. We then solve economic dispatch with all options of generators that can supply the load. Meaning, $\forall k:\;m\leq k\leq m_{\max}$, we set $y_{j\leq k}=1$ and $y_{j>k}=0$ and solve the following problem:
$\min\quad$ $\sum_{j=1}^{k}\left(a_{j}P_{j}^{2}+b_{j}P_{j}+c_{j}+K_{j}(1-u_{j}^{0})\right)$ (5a)
s.t. $P_{\min,j}\leq P_{j}\leq P_{\max,j}$ (5b)
$\sum_{j=1}^{k}P_{j}\geq D.$ (5c)
Eqs. (5a), (5b), and (5c) are the analogues of equations (2a), (2b), and (2c) with the binary variables $u_{i}$ set to $1$ or removed, because we have already decided which generators to operate. Eqs. (2d) and (2e) don’t have analogues, because they only depend on $u_{i}$ (not $P_{i}$) and we selected $k$ so that they hold. These $m_{max}-m+1$ problems are completely independent and can be solved in parallel, though Eq. 5 solves much faster than Eq. 2. Then we take the solution which had the lowest objective over all values of $m\leq k\leq m_{\max}$.
Algorithm 1 summarizes the workflow for this section.
Algorithm 1 Runtime constrained RRUC.
1:for $i=1$ to number_of_periods do
2:  Update rolling 24 hour generator states (use them in steps 3 and 4).
3:  Find must run generators (those whose minimum on time has not been reached).
4:  Find can’t run generators (those who have exceeded the maximum allowed daily starts).
5:  Solve Eq. 2 with $u_{i}\in\{0,1\}$ relaxed to $0\leq y_{i}\leq 1$.
6:  for Different sets of discretionary generators sufficient to meet the constraints (using Eq. 3 and Eq. 4) do
7:   Solve economic dispatch Eq. 5 with all must run generators and this discretionary generator set.
8:   if Objective is lowest to far then
9:    Update best solution and objective.
10:   end if
11:  end for
12:  return Best solution and best objective.
13:end for
Eq. 2 indicates that the problem has a time decoupled objective with time coupled constraints. This approach has a potential downside of selecting a different set of generating units compared to a fully coupled approach, because the future payoffs and penalties are not accounted for at a given simulation step. It has the advantage of simplifying the computation by reducing the number of required variables.
The results of this algorithm would not be directly comparable to fully coupled approach on computational time. Since we have shown our algorithm scales sub-quadratically [27], we believe it is possible to proceed with this approach and expand it in the future to include a time-coupled objective. In this work, the runtime constraints are still time-coupled and the system would be forced to converge while fully accounting for minimum runtime or the number of starts.
Unlike [27], we do not compare RRUC results to MIP results. The large system sizes, combined with new constraints, make MIP prohibitively computationally intensive for the available hardware (Apple M3 Pro, 18GB RAM). Instead we focus specifically on RRUC and adopt a new metric to assess the efficiency of the algorithm: percentage increase in the objective function and percentage increase in runtime as a function of increase in the system size. If the objective function and runtime do not explode as the system size increases. The algorithm maintains its scaling properties and is efficient.
III-B Simulation and Results
The system was simulated as follows. The historical PJM peak was recorded on June 23, 2025 at 5 pm, with total load of 160.2 GW  [24]. We apply this value to scale the system using a 20% cold reserve margin for resource adequacy, resulting in the system size of approximately 192 GW. Our selection of generating units is designed to represent the typical PJM generators plus the additional small-scale generators which we expect to be prevalent in the future as we explain in Section I. It includes 42 real PJM generators and amounts to 9047.9 MW capacity for the entire system, which we will refer to as $P_{\max}\doteq\sum_{i}P_{\max,i}$.
We multiply this generator set by 22 to pass 192 GW, using a random seed to rescale each generator in order to avoid the symmetry effects [19]. Non-identical copies of the original 42 generators have a shift in capacity parameters of $\pm 10\%$ and a shift in daily starts or runtime hour parameters of $\pm 1$. For various system sizes, we further use multiples of this system.
We solve the runtime constrained problem over 2304 time periods at 5 minute intervals (8 days). This covers the peak and adjacent days of 6/17/2025-6/24/2025. The latter four days of this period was the interval in which PJM experienced the highest change of load in 2025, with the night low of 78.6 GW and the day peak of 160.2 GW, with intraday changes of 55-64 GW. We start the simulation four days before the peak, so that our 72 hour forecasts have time to impact our unit commitments before we hit the peak. The first day’s results are used as a warm-up for the system (so that it is not starting with all units off) and we exclude them from our cost function. We assume that the standard deviation of the demand forecast is $1\text{GW}$ for the full PJM system (with $924$ generators) and scaled accordingly for larger systems.
To incorporate the runtime constraints, we augmented PJM generator data with runtime parameters summarized in Table I. PJM data had no clear assignment of units to gas or coal fuel, so we manually classified the units using the following rules: No unit larger than 1000 MW was considered gas because, according to EIA 860 [8], there were no operable gas units of this size in PJM territory. Coal units were identified as having a minimum runtime of 24 hours, unless they had the starting price in the supply curve above USD 10/MWh. The cost-based fuel assignment relies on the assumption that coal units have lower operating costs than gas units, although coal prices can be above gas prices on equivalent energy content [9].
Cold, warm, and hot start costs are available from PJM [24]. Those may be subject to strategic bidding, so it was helpful to validate them against a different system, such as the EU [3]. By normalizing warm and cold start costs by hot start cost, we obtain relative costs for coal and gas units that are broadly consistent: cold start is typically 1.4–2.3× the cost of hot start, and warm start is 1.1–1.6× the cost of hot start.
TABLE I: Runtime Assumptions and Data Sources
Parameter Source Coal Gas
Min runtime [min]  [24] 240-1440 0-1440
Max daily starts  [24] 1-3 1-24
Hot start [min]  [3, 13] 60-240 25-120
Warm start [min]  [3] 120-480 60-240
Cold start [min]  [3] 360-720 120-300
Relative cost of start, warm/hot  [24, 3] 1.19-1.42 1.12-1.58
Relative cost of start, cold/hot  [24, 3] 1.74-1.93 1.35-2.25
Figure 2: Runtime and objective function value for RRUC algorithm including runtime constraints with 462-14784 generators. With $n$ as the number of generators, the runtime scales as $\approx O(n^{3/2})$. This is believed to be ideal in the general case [20]. The objective function per generator increases less than 5% for each doubling of the system size.
We present the results of the simulation in Fig. 2 using the same software and hardware as before. Since the reference MIP solver is too slow to be used as a benchmark for RRUC, in this study we are mainly interested in analyzing the scaling and performance of RRUC. The runtime and the objective function increase slightly faster than linearly with the number of generators. Specifically, the runtime increases by less than $3\times$ for every doubling of the system size. The $\approx O(n^{3/2})$ scaling matches best case limits, due to a bottleneck from the linear solver [20]. For the objective function, we rely on the per unit measure, or the average total cost. The per-unit objective function increases less than 5% for every doubling of the system. An ideal solver’s objective per unit would not increase with the system size. This indicates that larger systems do not result in significant diseconomies of scale and the algorithm remains cost efficient.
IV Unit Commitment for Ramp and Sub Hourly Intervals
IV-A Mathematical Formulation
As the next step, we extend the problem to not only larger and more varied thermal power systems, but also systems that have non-thermal units with intermittent characteristics. To operate such systems, we must account for system behavior at a higher resolution, which requires a sub hour granularity of unit commitment. As discussed in Section I, high temporal resolution involves another important aspect of generator behavior, ramp.
There is a consensus among unit commitment studies that ramp is important in several ways [6, 23, 21]. First, it presents a constraint on how quickly a unit may be expected to enter operation. Second, a unit that has been connected to the system but has not reached $P_{\min,i}$ still produces a certain volume of active power. That active power enters the grid and changes the net demand that is left for other units. Accurately accounting for ramp output results in different objective function values and different sets of committed units.
Figure 3: A typical production cycle for generator $i$. When the generator is on, it can change its production power between $P_{\min,i}$ and $P_{\max,i}$, provided it does not exceed its ramping power. It can also turn off if it is close enough to $P_{\min,i}$ compared to its ability to ramp down. The only other decision can be made when the generator is off, to prepare for ramping up. During all other periods, the generator is uncontrollable.
To address the ramp constraints in the model, we use a different set of variables. Instead of the original $u_{i}$ that indicated that a generator is on or off, we allow the generator to be off, ramping up, on, or ramping down. For the model purposes, we assume that a generator can only change its status in one direction. If a generator is ramping up, it is assigned a variable $v_{i}=1$. After the required ramp time its status will switch to on. If a generator is ramping down, it is assigned a variable $w_{i}=1$. It will eventually switch off and enter the set of generators that can start. Fig. 3 shows a typical production cycle of a generator and the allowed transitions. Similarly to Section III, we use preprocessing to account for the accumulated runtime and ramping periods. The information about the duration of on, off, or ramping is stored for each generator and used as an optimization parameter.
The mathematical model used for ramp constrained sub hourly unit commitment is presented in Eq. 6.
$\min\quad$ $\sum_{i\in G_{m}}\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)$
$+\sum_{i\in G_{d}}\left[\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)u_{i}+K_{i}(w_{i}+v_{i})^{2}\right]$
$+\beta\sum_{i\in G_{d}}v_{i}\left(a_{i}P_{typ,i}+b_{i}+c_{i}/P_{typ,i}\right)$ (6a)
s.t. $\max(P_{\min,i},P_{t-1,i}-r_{d,i})\leq P_{i}$
$\leq\min(P_{\max,i},P_{t-1,i}+r_{u,i})$ (6b)
$u_{i}\in\{0,1\}\;v_{i}\in\{0,1\}\;w_{i}\in\{0,1\}$
$u_{i}+v_{i}+w_{i}\leq 1$ (6c)
$\sum_{i\in G_{m}}P_{i}+\sum_{i\in G_{d}}\left[u_{i}P_{i}+w_{i}(P_{\min,i}-r_{d,i})\right]\geq D-S_{r},$ (6d)
$\sum_{i\in G_{m}}P_{\max,i}+\sum_{i\in G_{d}}(u_{i}+v_{i})P_{\max,i}$
$\geq{D}_{\max,72}+3\sigma_{D}-S_{\max,r},$ (6e)
$\sum_{i\in G_{m}}P_{\min,i}+\sum_{i\in G_{d}}(u_{i}+v_{i})P_{\min,i}$
$\leq{D}_{\min,72}-3\sigma_{D}-S_{\min,r}.$ (6f)
The decision variables are:
  • $P_{i}\in[P_{\min,i},P_{\max,i}]$ for $i\in\mathcal{G}_{m}$: power output of must-run generator $i$
  • $P_{i}\in[P_{\min,i},P_{\max,i}]$ for $i\in\mathcal{G}_{d}$: power output of discretionary generator $i$
  • $v_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$ when the generator was previously off. 0 represents continuing to stay off, $1$ represents starting to ramp up.
  • $w_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$ when the generator was previously on. 0 represents continuing to stay on, $1$ represents starting to ramp down.
  • $u_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$: commitment status of discretionary generator $i$. We eliminate this variable in practice because it can be deduced from the previous status, $v_{i}$ and $w_{i}$.
The parameters supplied to the optimization solver are:
  • $a_{i},b_{i},c_{i}$: quadratic cost coefficients usually resulting from fuel costs or strategic bidding behavior.
  • $P_{\max,i},P_{\min,i}$: maximum capacity and minimum stable output, usually determined by physical characteristics of the units.
  • $D$, $\sigma_{D}$: demand volume and the possible demand deviations in the robust model formulation.
  • $K_{i}$: commitment change penalty for generator $i$. The true costs of startup and shutdown are usually not the same. Averaging the startup and shutdown costs prevents market distortions and ensures that there are no artificial incentives to run.
  • $r_{d,i}$ and $r_{u,i}$, generator $i$’s limits for ramp down or up in a given time period.
  • $\beta$ parameter we set to $0.1$. The corresponding term in the objective pushes us to choose more efficient generators for meeting the contingencies.
  • $P_{typ,i}$, a typical production level of a generator during a peak event, which we define as $(4*P_{\max,i}+P_{\min,i})/5$.
  • $P_{t-1,i}$ - production level for the generator at the previous time period. All variables are assumed to refer to the current time, unless otherwise noted.
  • $S_{r}$ - supply produced by generators that are already ramping before the solve (they cannot be changed). For generators ramping up, they are assumed to generate nothing for some number of periods, and then increase linearly to $P_{\min,i}$ in some number of time periods specific to the generator. Generators ramping down are assumed to decrease linearly from $P_{\min,i}$ in some number of time periods specific to the generator.
  • $S_{\max,r}$ total capacity of all generators ramping up (ramping down generators should be off by the time we reach the contingency)
  • $S_{\min,r}$ minimum operational capacity of all generators ramping up
  • ${D}_{\min,72}$ and ${D}_{\max,72}$: minimum and maximum volumes of predicted demand in next 72 hours.
  • $\mathcal{G}_{m}$: the set of must run generators.
  • $\mathcal{G}_{d}$: the set of discretionary generators.
  • $R=\max\{\max_{i\in\mathcal{G}_{m}}P_{\max,i},\max_{i\in\mathcal{G}_{d}}P_{\max,i}\}$: reserve margin (changed to only include generators that can run).
Algorithm 2 summarizes the workflow for this section.
Algorithm 2 Runtime and ramp constrained RRUC.
1:for $i=1$ to number_of_periods do
2:  Update rolling 24 hour generator states (use them in steps 3-8).
3:  Increment the number of ramping periods for all ramping generators.
4:  Switch generators that have finished preparation to ramping up.
5:  Switch generators that have finished ramping up to on.
6:  Switch generators that have finished ramping down to off.
7:  Find must run generators (those whose minimum on time has not been reached or are producing too far above $P_{\min,i}$ to shut off).
8:  Find can’t start generators (those who have exceeded the maximum allowed daily starts or are already ramping).
9:  Solve Eq. 6 with $u_{i}\in\{0,1\}$ relaxed to $0\leq y_{i}\leq 1$.
10:  for Different sets of discretionary generators sufficient to meet the constraints (using Eq. 3 and Eq. 4) do
11:   Solve economic dispatch Eq. 5 with all must run generators and this discretionary generator set.
12:   if Objective is lowest so far then
13:    Update best solution and objective.
14:   end if
15:  end for
16:  return Best solution and best objective.
17:end for
IV-B Simulation and Results
Much of the data used in this Section has already been presented in Section III for runtime constraints. The additional assumptions about equipment characteristics are presented in Table II. While ramp characteristics usually come in MW/min, we use the % of $P_{\max,i}$ per minute as the original input into the simulation. The percentage values are compiled from several sources, including the academic reports from Deutsches Institut für Wirtschaftsforschung [3] and manufacturer catalog of GE Vernova [13]. The provided numbers, although summarising different sources, may still present the upper bound of ramp rates, as comprehensive survey of large thermal units suggest that in real systems power plants do not always deliver the nominal ramp rates specified for equipment [25].
Not all the available generating units could be accurately matched to the PJM set of generators on $P_{\min,i}$ and $P_{\max,i}$. Therefore we used approximations based on size and the assumed fuel used by a turbine (gas or coal steam). Larger turbines were assigned lower ramp values, as were coal steam turbines. Once the ramp percentages were assigned to specific generating units, the $P_{\max,i}$ of those generators was multiplied by the ramp rate to get the MW/min value, which was later used in the simulation.
TABLE II: Ramp Assumptions and Data Sources
Parameter Source Coal Gas
Ramp up rate [%/min]  [3],  [13] 1-6 2-12
Ramp down rate [%/min]  [3] 5 15
Figure 4: Runtime and objective function value for RRUC algorithm with runtime constraints and ramping constraints for 462-14784 generators. The runtime scales as $\approx O(n^{4/3})$, where $n$ is the number of generators. The objective function per generator increases less than 2% for each doubling of the system size.
The runtime and objective results are summarized in Fig. 4 and are comparable to those found in Section III using the same software and hardware. The runtime scaled roughly as $n^{4/3}$, where $n$ is the number of generators. This shows that the resulting linear systems are less complicated than those arising from a typical optimization problem ($n^{3/2}$) and that our optimization solver is close to ideal. It is well below the exponential cost of MIP solvers. The average cost of the system increased by at most 2% for each doubling of the system size. So adding extra generators does not result in significant scale cost effects.
V Unit Commitment With Non-constant Ramps
Finally, we tested the suggested RRUC methodology on systems with non-constant ramp rates. The use of constant ramp rates is common in research, but they represent a simplified view of generator constraints. The ramp rates are not identical for ramping up and ramping down[3]. Ramping rates are also not constant during the ramping process, later stages of ramping are different from earlier stages[6], [31]. To account for the different ramp rates, we introduce a model extension that applies different ramp rates during different stages of generator operation.
The results were almost indistinguishable from Section IV. This is likely due to the fact that our solver is highly averse to turning units off or on, because it pays a price for it, and the generation payoff only happens in later time periods. It will only switch units on or off if it is required to meet the constraints, which happens very rarely. This motivates future work where we define our optimization objective over all time periods, instead of greedily.
VI Conclusion and Further Work
We proposed an algorithm to solve unit commitment with mechanical and ramping constraints by relaxing the integer constraints and then rounding solutions so that constraints are satisfied. Our algorithm shows $\approx O(n^{4/3})-O(n^{3/2})$ time scaling in the number of generators. It retains accuracy nearly perfectly, as the problem size increases.
For a system with runtime constraints, a $2\times$ increase in the number of units results in less than $3\times$ increase in the computational time. Further, the average value of the objective function increased by less than $1.05\times$ for each $2\times$ increase in the system size, indicating that there are no diseconomies of scale associated with the increase in system size using the suggested algorithm.
When ramping constraints were added to the problem, runtime scaling was even closer to linear. The objective (cost) function per generator increased by less than 2% for each $2\times$ increase in the system size.
A switch from two ramping rates, one for ramping up and one for ramping down, to four rates for early ramping up, later ramping up, early ramping down and late ramping down, had almost no effect on the results. This may indicate that the algorithm is robust to changes in the ramp rate inputs, provided the allowed rates are fast enough.
Further research would be required to assess the performance of the proposed methods when applied to a fully temporally coupled objective function and constraints. Further, it would be valuable to identify if the suggested algorithm demonstrates good convergence and scaling in the presence of transmission constraints.
Acknowledgements
The authors thank Nicholson Koukpaizan for contributing to an improved manuscript with his thorough technical review.

### v2

@@
-A Fast Relax-and-Round Approach to Unit Commitment with Sub Hourly Mechanical and Ramp Constraints
+Fast Relax-and-Round Unit Commitment with Sub-hourly Mechanical and Ramp Constraints
 Abstract
@@
 Running 1000 unit commitment scenarios for a 150-unit system remains substantially faster than running one scenario for a 1000-unit system.
-Similarly, running 72 time decoupled unit commitment scenarios is faster than running one time coupled scenario for a three days outlook.
+Similarly, running 72 time decoupled unit commitment scenarios is faster than running one time-coupled scenario for a three days outlook.
 Second, existing unit commitment algorithms rely on mixed integer programs (MIPs) [2].
@@
 To achieve this, we first review the basic UC formulation from [27] and then augment it with supplementary constraints usually found in literature [6, 19].
-Our approach only accounts for the costs of generation in the current period, but includes intertemporal runtime constraints and, later, ramping constraints.
+Our approach only accounts for the actual costs of generation in the current period, but is biased towards turning more efficient units on.
+It includes intertemporal runtime constraints and, later, ramping constraints.
 From that perspective, the algorithm is partially temporally decoupled and cannot be directly compared to fully coupled UC algorithms.
@@
 Section IV extends the model to include ramping constraints.
-Section V briefly discusses results with a different ramping formulation.
+Section V introduces a different ramping formulation and compares its results to Section IV.
 Section VI summarizes the results and suggests further research directions.
@@
   • $D$: current predicted demand.
-  • $\sigma_{D}$: standard deviation of demand forecast
+  • $\sigma_{D}$: standard deviation (STD) of demand forecast
   • $R=\max\{P_{\max,i}\}$: reserve margin.
@@
 The constraint Eq.
-1d ensures that if our largest generator goes offline we will be able to supply demand if it is within 3 standard deviations of our prediction.
+1d ensures that if our largest generator goes offline we will be able to supply demand if it is within $3\sigma_{D}$ of our prediction.
 This ensures we can still supply demand if any one generator goes out and our predicted demand is higher than expected.
@@
 $\sum_{i\in G_{m}}P_{\max,i}+\sum_{i\in G_{d}}u_{i}P_{\max,i}\geq{D}_{\max,72}+3\sigma_{D}+R,$ (2d)
-$\sum_{i\in G_{m}}P_{\min,i}+\sum_{i\in G_{d}}u_{i}P_{\min,i}\leq{D}_{\min,72}-3\sigma_{D}.$ (2e)
+$\sum_{i\in G_{m}}P_{\min,i}+\sum_{i\in G_{d}}u_{i}P_{\min,i}\leq{D}_{\min,72}-\sigma_{D}.$ (2e)
+Note that at the lower end, we plan for a smaller contingency ($1\sigma_{D}$) than on the higher end ($3\sigma_{D}$).
+This is meant to reflect the asymmetric nature of load forecasts.
+It is far more likely that load will be much higher than predicted (corresponding to some large unexpected load), compared to much lower (corresponding to base load being removed).
+This also helps provide additional buffers in the event that more than 1 generator goes out, or uncontrollable generation is much lower than expected.
 The parameters supplied to the optimization solver are:
@@
 4:
-$\sum_{i\in\mathcal{G}_{m}}P_{\min,i}+\sum_{\begin{subarray}{c}1\leqslant i\leqslant\tilde{m}_{\max}\\ i\in\mathcal{G}_{d}\end{subarray}}P_{\min,i}\;\geq\;D_{\min,72}+3\sigma_{D},$ (4)
+$\sum_{i\in\mathcal{G}_{m}}P_{\min,i}+\sum_{\begin{subarray}{c}1\leqslant i\leqslant\tilde{m}_{\max}\\ i\in\mathcal{G}_{d}\end{subarray}}P_{\min,i}\;\geq\;D_{\min,72}-\sigma_{D},$ (4)
 with $m_{\max}\doteq|\mathcal{G}_{m}|+\tilde{m}_{\max}$.
@@
 Eq.
-2 indicates that the problem has a time decoupled objective with time coupled constraints.
+2 indicates that the problem has a weakly time-coupled objective with strongly time-coupled constraints.
 This approach has a potential downside of selecting a different set of generating units compared to a fully coupled approach, because the future payoffs and penalties are not accounted for at a given simulation step.
@@
 The results of this algorithm would not be directly comparable to fully coupled approach on computational time.
-Since we have shown our algorithm scales sub-quadratically [27], we believe it is possible to proceed with this approach and expand it in the future to include a time-coupled objective.
+Since we have shown our algorithm scales sub-quadratically [27], we believe it is possible to proceed with this approach and expand it in the future to include a fully time-coupled objective.
 In this work, the runtime constraints are still time-coupled and the system would be forced to converge while fully accounting for minimum runtime or the number of starts.
@@
 The first day’s results are used as a warm-up for the system (so that it is not starting with all units off) and we exclude them from our cost function.
-We assume that the standard deviation of the demand forecast is $1\text{GW}$ for the full PJM system (with $924$ generators) and scaled accordingly for larger systems.
+We assume that the standard deviation of the demand forecast $\sigma_{D}=1\text{GW}$ for the full PJM system (with $924$ generators) and scaled accordingly for larger systems.
 To incorporate the runtime constraints, we augmented PJM generator data with runtime parameters summarized in Table I.
@@
 Figure 2: Runtime and objective function value for RRUC algorithm including runtime constraints with 462-14784 generators.
-With $n$ as the number of generators, the runtime scales as $\approx O(n^{3/2})$.
-This is believed to be ideal in the general case [20].
+With $n$ as the number of generators, the runtime scales as $\approx O(n^{1.37})$.
 The objective function per generator increases less than 5% for each doubling of the system size.
@@
 The runtime and the objective function increase slightly faster than linearly with the number of generators.
-Specifically, the runtime increases by less than $3\times$ for every doubling of the system size.
-The $\approx O(n^{3/2})$ scaling matches best case limits, due to a bottleneck from the linear solver [20].
-For the objective function, we rely on the per unit measure, or the average total cost.
+Specifically, the runtime increases on average about $2.6\times$ for every doubling of the system size.
+The $\approx O(n^{1.37})$ shows that the resulting systems are less complicated than those arising from a typical optimization problem which have $O(n^{1.5})$ scaling.
+For a general optimization problem, $O(n^{1.5})$ scaling is believed to be ideal, due to the bottleneck from the linear solver [20].
+For the objective function, we rely on the per unit measure, or the average total cost, because our demand is scaled accordingly.
 The per-unit objective function increases less than 5% for every doubling of the system.
+It is possible we could avoid the small loss of solution quality by increasing the run time.
 An ideal solver’s objective per unit would not increase with the system size.
+However, this model is just a stepping stone to the more realistic one in  Section IV, so we will not spend time optimizing it.
 This indicates that larger systems do not result in significant diseconomies of scale and the algorithm remains cost efficient.
@@
 IV-A Mathematical Formulation
-As the next step, we extend the problem to not only larger and more varied thermal power systems, but also systems that have non-thermal units with intermittent characteristics.
+We now extend the problem to larger and more varied thermal power systems, which have have non-thermal units with intermittent characteristics.
 To operate such systems, we must account for system behavior at a higher resolution, which requires a sub hour granularity of unit commitment.
@@
 Accurately accounting for ramp output results in different objective function values and different sets of committed units.
-Figure 3: A typical production cycle for generator $i$.
+Figure 3: A typical production cycle for generator $i$ in Section IV.
 When the generator is on, it can change its production power between $P_{\min,i}$ and $P_{\max,i}$, provided it does not exceed its ramping power.
@@
 To address the ramp constraints in the model, we use a different set of variables.
-Instead of the original $u_{i}$ that indicated that a generator is on or off, we allow the generator to be off, ramping up, on, or ramping down.
+Instead of the original $u_{i}$ that indicated that a generator is on or off, we allow the generator to be off, preparing (for ramp up), ramping up, on, or ramping down.
 For the model purposes, we assume that a generator can only change its status in one direction.
-If a generator is ramping up, it is assigned a variable $v_{i}=1$.
-After the required ramp time its status will switch to on.
+If a generator is ramping up or preparing, it is assigned a variable $v_{i}=1$.
+After the required ramp time in each stage, its status will switch to on.
 If a generator is ramping down, it is assigned a variable $w_{i}=1$.
@@
 3 shows a typical production cycle of a generator and the allowed transitions.
+Each generator can only be in one state in a given time period, so we round ramping times from the data to whole periods.
+We round up to ensure that our solution is for a problem more restrictive than the actual problem, which means it is necessarily feasible.
+This means a generator must occupy each state for at least one time period.
 Similarly to Section III, we use preprocessing to account for the accumulated runtime and ramping periods.
@@
 $\sum_{i\in G_{m}}P_{\min,i}+\sum_{i\in G_{d}}(u_{i}+v_{i})P_{\min,i}$
-$\leq{D}_{\min,72}-3\sigma_{D}-S_{\min,r}.$ (6f)
+$\leq{D}_{\min,72}-\sigma_{D}-S_{\min,r}.$ (6f)
 The decision variables are:
@@
   • $r_{d,i}$ and $r_{u,i}$, generator $i$’s limits for ramp down or up in a given time period.
-  • $\beta$ parameter we set to $0.1$. The corresponding term in the objective pushes us to choose more efficient generators for meeting the contingencies.
+  • $\beta$ parameter we set to $0.001$. The corresponding term in the objective pushes us to choose more efficient generators for meeting the contingencies.
   • $P_{typ,i}$, a typical production level of a generator during a peak event, which we define as $(4*P_{\max,i}+P_{\min,i})/5$.
@@
 Ramp down rate [%/min]  [3] 5 15
-Figure 4: Runtime and objective function value for RRUC algorithm with runtime constraints and ramping constraints for 462-14784 generators.
-The runtime scales as $\approx O(n^{4/3})$, where $n$ is the number of generators.
-The objective function per generator increases less than 2% for each doubling of the system size.
-The runtime and objective results are summarized in Fig.
-4 and are comparable to those found in Section III using the same software and hardware.
-The runtime scaled roughly as $n^{4/3}$, where $n$ is the number of generators.
-This shows that the resulting linear systems are less complicated than those arising from a typical optimization problem ($n^{3/2}$) and that our optimization solver is close to ideal.
+Figure 4: Runtime and objective function value for RRUC algorithm with runtime constraints and ramping constraints for 462-14784 generators with the Fig.
+3 profile.
+The objective function per generator fluctuates, but does not show an increasing or decreasing trend based on the system size.
+The runtime scales as $\approx O(n^{1.6})$, where $n$ is the number of generators.
+Fig.
+4 summarizes the runtime and objective results.
+These results are comparable to those found in Section III using the same software and hardware.
+The runtime scaled roughly as $n^{1.6}$, where $n$ is the number of generators.
+Ideal scaling is believed to be bounded below by $n^{1.5}$, due to the scaling of the underlying linear system solvers [20].
+This suggests that our optimization solver has close to ideal scaling.
 It is well below the exponential cost of MIP solvers.
-The average cost of the system increased by at most 2% for each doubling of the system size.
-So adding extra generators does not result in significant scale cost effects.
-V Unit Commitment With Non-constant Ramps
-Finally, we tested the suggested RRUC methodology on systems with non-constant ramp rates.
+The average cost of the system did not increase with the system size.
+So adding extra generators does not result in scale cost effects.
+V Unit Commitment With Smooth Ramp Up
+Finally, we tested the suggested RRUC methodology on systems with smooth ramp up.
 The use of constant ramp rates is common in research, but they represent a simplified view of generator constraints.
@@
 Ramping rates are also not constant during the ramping process, later stages of ramping are different from earlier stages[6], [31].
-To account for the different ramp rates, we introduce a model extension that applies different ramp rates during different stages of generator operation.
-The results were almost indistinguishable from Section IV.
-This is likely due to the fact that our solver is highly averse to turning units off or on, because it pays a price for it, and the generation payoff only happens in later time periods.
-It will only switch units on or off if it is required to meet the constraints, which happens very rarely.
-This motivates future work where we define our optimization objective over all time periods, instead of greedily.
+Further, there is no clear distinction in practice between the prepare and ramp up stages in Fig.
+3.
+They are both part of a deterministic shift from the generator being off to on.
+Figure 5: A typical production cycle for generator $i$ in Section V.
+When the generator is on, it can change its production power between $P_{\min,i}$ and $P_{\max,i}$, provided it does not exceed its ramping power.
+It can also turn off if it is close enough to $P_{\min,i}$ compared to its ability to ramp down.
+The only other decision can be made when the generator is off, start ramping up.
+During all other periods, the generator is uncontrollable.
+We introduce a model extension in Fig.
+5 that accounts for these discrepancies to create a more realistic model of generator states.
+We combine the prepare and ramp up stages into one stage where production increases quadratically from 0 to $P_{\min,i}$.
+This produces a more realistic, smoothed version of Fig.
+3.
+Due to rounding up to whole time periods, this has the added effect of potentially requiring one less period for ramping up, which is closer to the actual ramp up time.
+Figure 6: Runtime and objective function value for RRUC algorithm with runtime constraints and ramping constraints for 462-14784 generators, with the Fig.
+5 profile.
+The objective function per generator fluctuates, but does not show an increasing or decreasing trend based on the system size.
+The runtime scales as $\approx O(n^{1.6})$.
+Fig.
+6 shows results using the Fig.
+5 generator profile model.
+The results were similar to those of Section IV.
+There is no upward or downward trend in objective value per generator as the number of generators grows.
+Runtime continued to scale sub-quadratically, slightly slower than the optimal $O(n^{1.5})$.
+Figure 7: Number of generators in a given state for the 462 generator simulation.
+Ramp up includes the preparation stage for Section IV results.
+Note: Only every 66th time step is shown for readability.
+(Top) Total number of generators in the Section IV simulation in on, ramp up, and ramp down states.
+Where there are 0 generators in a given state, it is omitted on the log-scale.
+At almost all time steps, the majority of active generators are on and controllable.
+(Bottom) Difference in the number of generators in a given state between the simulation in Section IV (Fig.
+3 profile) and Section V (Fig.
+5 profile).
+At almost all time steps, there is no difference.
+The top part of Fig.
+7 shows that most time steps have no generators ramping up or down.
+In $1881$ out of $2016$ time steps ($93.3\%$), all generators were on or off for the Fig.
+3 generator model.
+In $1884$ out of $2016$ time steps ($93.5\%$), all generators were on or off for the Fig.
+5 generator model.
+Ramp up and ramp down bursts tend to be short and include many generators.
+The bottom part of Fig.
+7 shows the discrepancies between the solutions that both generator profiles produce.
+The differences in the generator states of the solutions remain large for the first $\approx 800$ time steps (equivalent to almost $3$ days), but level off.
+Small differences in the input can lead to very different but similar quality (similar cost) solutions, when aggregated over all time periods.
+This shows the non-convexity of the problem and the strength of the optimization solver, which is able to find different minima.
 VI Conclusion and Further Work
-We proposed an algorithm to solve unit commitment with mechanical and ramping constraints by relaxing the integer constraints and then rounding solutions so that constraints are satisfied.
-Our algorithm shows $\approx O(n^{4/3})-O(n^{3/2})$ time scaling in the number of generators.
-It retains accuracy nearly perfectly, as the problem size increases.
-For a system with runtime constraints, a $2\times$ increase in the number of units results in less than $3\times$ increase in the computational time.
-Further, the average value of the objective function increased by less than $1.05\times$ for each $2\times$ increase in the system size, indicating that there are no diseconomies of scale associated with the increase in system size using the suggested algorithm.
-When ramping constraints were added to the problem, runtime scaling was even closer to linear.
-The objective (cost) function per generator increased by less than 2% for each $2\times$ increase in the system size.
-A switch from two ramping rates, one for ramping up and one for ramping down, to four rates for early ramping up, later ramping up, early ramping down and late ramping down, had almost no effect on the results.
-This may indicate that the algorithm is robust to changes in the ramp rate inputs, provided the allowed rates are fast enough.
+We proposed an algorithm to solve unit commitment with mechanical and ramping constraints by relaxing the integer constraints and then rounding solutions to satisfy the constraints.
+Our algorithm shows between $\approx O(n^{1.6})$ time scaling in the number of generators.
+As the problem size increases, it retains accuracy nearly perfectly.
+For a system with only runtime constraints, a $2\times$ increase in the number of units results in about a $2.6\times$ increase in the computational time on average ($\approx O(n^{1.37})$ scaling).
+Further, the average value of the objective function increased by less than $1.05\times$ for each $2\times$ increase in the system size, indicating that there are some small diseconomies of scale associated with the increase in system size using the suggested algorithm.
+When ramping constraints were added to the problem, runtime scaling was close to ideal.
+For every $2\times$ increase in problem size, runtime increased about $3\times$ on average.
+The objective (cost) function per generator did not increase with the problem size.
+It is possible that being more restrictive is beneficial for a ramp constrained formulation for objective scaling.
+The demand at each time step is close to the one preceding it, so a more restricted solver with ramping constraints can restrict the search on better solutions.
+This means that even as the problem grows, at the cost of more runtime, the solver can pinpoint similar quality solutions.
+A switch to a smoother generator profile did not change the average results drastically.
+Objectives and their perfect scaling were similar.
+runtimes and their scaling were slightly better.
+For every $2\times$ increase in the problem size, the runtime increased less than $3\times$ on average.
+However, depending on the generator profile, certain time periods have substantially different units in different states.
+This indicates that even slightly different ramping profiles can lead to very different unit commitments, even if their objectives are similar.
+This highlights the non-convexity of the problem and the importance of accurate knowledge about generator profiles.
 Further research would be required to assess the performance of the proposed methods when applied to a fully temporally coupled objective function and constraints.

### v3

@@
 Generator EIA data shows they are becoming smaller.
-For units in operation in 2024, excluding Alaska and Hawaii, average nameplate capacity for a generator built in 1974 was 205.22 MW.
+For US units in operation in 2024, excluding Alaska and Hawaii, average nameplate capacity for a generator built in 1974 was 205.22 MW.
 For a generator built in 2024 it was 49.5 MW [8].
@@
   • $D$: current predicted demand.
-  • $\sigma_{D}$: standard deviation (STD) of demand forecast
+  • $\sigma_{D}$: standard deviation of demand forecast
   • $R=\max\{P_{\max,i}\}$: reserve margin.
@@
 The resulting optimization problem is:
-$\min\quad$ $\sum_{i\in G_{m}}\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)$
-$+\sum_{i\in G_{d}}\left[\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)u_{i}+K_{i}(u_{i}-u_{i}^{0})^{2}\right]$ (2a)
+$\min\quad$ $\sum_{i\in G_{m}}\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)+$ (2a)
+$\sum_{i\in G_{d}}\left[\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)u_{i}+K_{i}(u_{i}-u_{t-1,i})^{2}\right]$
 s.t. $P_{\min,i}\leq P_{i}\leq P_{\max,i},\;u_{i}\in\{0,1\}$ (2b)
@@
 $\sum_{i\in G_{m}}P_{\min,i}+\sum_{i\in G_{d}}u_{i}P_{\min,i}\leq{D}_{\min,72}-\sigma_{D}.$ (2e)
-Note that at the lower end, we plan for a smaller contingency ($1\sigma_{D}$) than on the higher end ($3\sigma_{D}$).
-This is meant to reflect the asymmetric nature of load forecasts.
-It is far more likely that load will be much higher than predicted (corresponding to some large unexpected load), compared to much lower (corresponding to base load being removed).
-This also helps provide additional buffers in the event that more than 1 generator goes out, or uncontrollable generation is much lower than expected.
 The parameters supplied to the optimization solver are:
   • $a_{i},b_{i},c_{i}$: quadratic cost coefficients usually resulting from fuel costs or strategic bidding behavior.
-  • $P_{\max,i},P_{\min,i}$: maximum capacity and minimum stable output, usually determined by mechanical characteristics of the units.
+  • $P_{\max,i},P_{\min,i}$: maximum capacity and minimum stable output, determined by a unit’s physical characteristics
   • $D$, $\sigma_{D}$: demand volume and the possible demand deviations in the robust model formulation.
-  • $K_{i}$: commitment change penalty for generator $i$.
-The true costs of startup and shutdown are usually not the same.
-Averaging the startup and shutdown costs prevents market distortions and ensures that there are no artificial incentives to run.
-  • $u_{i}^{0}\in\{0,1\}$: previous commitment status of generator $i$
+  • $K_{i}$: commitment change penalty for generator ii [^2] The true startup and shutdown costs are rarely the same.
+Averaging them prevents market distortion and removes artificial incentives to turn on or off..
+  • $u_{t-1,i}\in\{0,1\}$: previous commitment status of unit $i$
   • ${D}_{\min,72}$ and ${D}_{\max,72}$: minimum and maximum volumes of predicted demand in next 72 hours.
@@
 Meaning, $\forall k:\;m\leq k\leq m_{\max}$, we set $y_{j\leq k}=1$ and $y_{j>k}=0$ and solve the following problem:
-$\min\quad$ $\sum_{j=1}^{k}\left(a_{j}P_{j}^{2}+b_{j}P_{j}+c_{j}+K_{j}(1-u_{j}^{0})\right)$ (5a)
+$\min\quad$ $\sum_{j=1}^{k}\left(a_{j}P_{j}^{2}+b_{j}P_{j}+c_{j}+K_{j}(u_{j}-u_{t-1,j})^{2}\right)$ (5a)
 s.t. $P_{\min,j}\leq P_{j}\leq P_{\max,j}$ (5b)
@@
 Algorithm 1 Runtime constrained RRUC.
-1:for $i=1$ to number_of_periods do
-2:  Update rolling 24 hour generator states (use them in steps 3 and 4).
-3:  Find must run generators (those whose minimum on time has not been reached).
-4:  Find can’t run generators (those who have exceeded the maximum allowed daily starts).
+1:for $i=1$ to $\mathrm{number\_of\_periods}$ do
+2:  Update rolling 24-hour generator states (use them in steps 3 and 4).
+3:  Find must run units (with minimum on time not yet reached or with $P_{t-1,i}-P_{min,i}$ too large to shut off)
+4:  Find can’t start generators (those that have exceeded maximum allowed daily starts or are already ramping)
 5:  Solve Eq. 2 with $u_{i}\in\{0,1\}$ relaxed to $0\leq y_{i}\leq 1$.
-6:  for Different sets of discretionary generators sufficient to meet the constraints (using Eq. 3 and Eq. 4) do
-7:   Solve economic dispatch Eq. 5 with all must run generators and this discretionary generator set.
-8:   if Objective is lowest to far then
-9:    Update best solution and objective.
-10:   end if
-11:  end for
-12:  return Best solution and best objective.
-13:end for
+6:  Order units in $\mathcal{G}_{d}$ into sets descending by $y_{i}$
+7:  for each $\mathcal{G}_{d}$ that meets (Eq. 2b)-(Eq. 2e) do
+8:   Solve economic dispatch Eq. 5 with $\mathcal{G}_{m},\mathcal{G}_{d}$ .
+9:   if objective is lowest so far then
+10:     Update best solution and objective.
+11:   end if
+12:  end for
+13:  save Best solution and best objective.
+14:end for
 Eq.
@@
 It has the advantage of simplifying the computation by reducing the number of required variables.
-The results of this algorithm would not be directly comparable to fully coupled approach on computational time.
+Results from Algorithm 1 are not be directly comparable to fully coupled approach on computational time.
 Since we have shown our algorithm scales sub-quadratically [27], we believe it is possible to proceed with this approach and expand it in the future to include a fully time-coupled objective.
-In this work, the runtime constraints are still time-coupled and the system would be forced to converge while fully accounting for minimum runtime or the number of starts.
+In this work, the runtime constraints are still time-coupled and the system is forced to converge while fully accounting for minimum runtime or the number of starts.
 Unlike [27], we do not compare RRUC results to MIP results.
@@
 Instead we focus specifically on RRUC and adopt a new metric to assess the efficiency of the algorithm: percentage increase in the objective function and percentage increase in runtime as a function of increase in the system size.
-If the objective function and runtime do not explode as the system size increases.
-The algorithm maintains its scaling properties and is efficient.
+If the objective function and runtime do not explode as the system size increases, the algorithm maintains its scaling properties and is efficient.
 III-B Simulation and Results
@@
 We multiply this generator set by 22 to pass 192 GW, using a random seed to rescale each generator in order to avoid the symmetry effects [19].
-Non-identical copies of the original 42 generators have a shift in capacity parameters of $\pm 10\%$ and a shift in daily starts or runtime hour parameters of $\pm 1$.
+The noisy copies of the original 42 generators have a shift in capacity parameters of $\pm 10\%$ and a shift in daily starts or runtime hour parameters of $\pm 1$.
 For various system sizes, we further use multiples of this system.
@@
 The $\approx O(n^{1.37})$ shows that the resulting systems are less complicated than those arising from a typical optimization problem which have $O(n^{1.5})$ scaling.
-For a general optimization problem, $O(n^{1.5})$ scaling is believed to be ideal, due to the bottleneck from the linear solver [20].
+For a general optimization problem, $O(n^{1.5})$ scaling is believed to be ideal, due to the linear solver bottleneck from the  [20].
 For the objective function, we rely on the per unit measure, or the average total cost, because our demand is scaled accordingly.
@@
 To operate such systems, we must account for system behavior at a higher resolution, which requires a sub hour granularity of unit commitment.
-As discussed in Section I, high temporal resolution involves another important aspect of generator behavior, ramp.
+Ramping is another important aspect of generator behavior at high temporal resolutions.
 There is a consensus among unit commitment studies that ramp is important in several ways [6, 23, 21].
@@
 After the required ramp time in each stage, its status will switch to on.
-If a generator is ramping down, it is assigned a variable $w_{i}=1$.
-It will eventually switch off and enter the set of generators that can start.
+If a generator is ramping down, it will eventually switch off and enter the set of generators that can start.
 Fig.
@@
 $\min\quad$ $\sum_{i\in G_{m}}\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)$
-$+\sum_{i\in G_{d}}\left[\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)u_{i}+K_{i}(w_{i}+v_{i})^{2}\right]$
+$+\sum_{i\in G_{d}}\left[\left(a_{i}P_{i}^{2}+b_{i}P_{i}+c_{i}\right)u_{i}+K_{i}(1-u_{i}+v_{i})^{2}\right]$
 $+\beta\sum_{i\in G_{d}}v_{i}\left(a_{i}P_{typ,i}+b_{i}+c_{i}/P_{typ,i}\right)$ (6a)
@@
 $\leq\min(P_{\max,i},P_{t-1,i}+r_{u,i})$ (6b)
-$u_{i}\in\{0,1\}\;v_{i}\in\{0,1\}\;w_{i}\in\{0,1\}$
-$u_{i}+v_{i}+w_{i}\leq 1$ (6c)
-$\sum_{i\in G_{m}}P_{i}+\sum_{i\in G_{d}}\left[u_{i}P_{i}+w_{i}(P_{\min,i}-r_{d,i})\right]\geq D-S_{r},$ (6d)
+$u_{i}\in\{0,1\},\;v_{i}\in\{0,1\},\;u_{i}+v_{i}\leq 1$ (6c)
+$\sum_{i\in G_{m}}P_{i}+\sum_{i\in G_{d}}\left[u_{i}P_{i}+(1-u_{i})(P_{\min,i}-r_{d,i})\right]$
+$\geq D-S_{r},$ (6d)
 $\sum_{i\in G_{m}}P_{\max,i}+\sum_{i\in G_{d}}(u_{i}+v_{i})P_{\max,i}$
@@
   • $P_{i}\in[P_{\min,i},P_{\max,i}]$ for $i\in\mathcal{G}_{d}$: power output of discretionary generator $i$
-  • $v_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$ when the generator was previously off. 0 represents continuing to stay off, $1$ represents starting to ramp up.
-  • $w_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$ when the generator was previously on. 0 represents continuing to stay on, $1$ represents starting to ramp down.
-  • $u_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$: commitment status of discretionary generator $i$.
-We eliminate this variable in practice because it can be deduced from the previous status, $v_{i}$ and $w_{i}$.
+  • $u_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$: commitment status of discretionary generator $i$. $u_{i}=0$ is off and $u_{i}=1$ is on.
+  • $v_{i}\in\{0,1\}$ for $i\in\mathcal{G}_{d}$ when $u_{i}$ was previously 1. $v_{i}=0$ represents staying off and $v_{i}=1$ is starting to ramp up.
 The parameters supplied to the optimization solver are:
   • $a_{i},b_{i},c_{i}$: quadratic cost coefficients usually resulting from fuel costs or strategic bidding behavior.
-  • $P_{\max,i},P_{\min,i}$: maximum capacity and minimum stable output, usually determined by physical characteristics of the units.
+  • $P_{\max,i},P_{\min,i}$: maximum capacity and minimum stable output, determined by a unit’s physical characteristics.
   • $D$, $\sigma_{D}$: demand volume and the possible demand deviations in the robust model formulation.
   • $K_{i}$: commitment change penalty for generator $i$.
-The true costs of startup and shutdown are usually not the same.
-Averaging the startup and shutdown costs prevents market distortions and ensures that there are no artificial incentives to run.
-  • $r_{d,i}$ and $r_{u,i}$, generator $i$’s limits for ramp down or up in a given time period.
-  • $\beta$ parameter we set to $0.001$. The corresponding term in the objective pushes us to choose more efficient generators for meeting the contingencies.
-  • $P_{typ,i}$, a typical production level of a generator during a peak event, which we define as $(4*P_{\max,i}+P_{\min,i})/5$.
+  • $r_{d,i}$ and $r_{u,i}$: generator $i$’s limits for ramp down normalized so that $\Delta t=1$.
+  • $\beta$: bias factor for selecting more efficient generators to meet the contingencies. In our model we set it to $0.001$.
+  • $P_{typ,i}$, a typical production level of a generator during a peak event, which we define as $(4P_{\max,i}+P_{\min,i})/5$.
   • $P_{t-1,i}$ - production level for the generator at the previous time period. All variables are assumed to refer to the current time, unless otherwise noted.
-  • $S_{r}$ - supply produced by generators that are already ramping before the solve (they cannot be changed).
-For generators ramping up, they are assumed to generate nothing for some number of periods, and then increase linearly to $P_{\min,i}$ in some number of time periods specific to the generator.
-Generators ramping down are assumed to decrease linearly from $P_{\min,i}$ in some number of time periods specific to the generator.
-  • $S_{\max,r}$ total capacity of all generators ramping up (ramping down generators should be off by the time we reach the contingency)
-  • $S_{\min,r}$ minimum operational capacity of all generators ramping up
+  • $S_{r}$ - supply produced by generators that are already ramping before the solve (they cannot be changed) [^3] For generators ramping up, they are assumed to generate nothing for some number of periods, and then increase linearly to $P_{\min,i}$ in some number of time periods specific to the generator.
+Generators ramping down are assumed to decrease linearly from $P_{\min,i}$ in some number of time periods specific to the generator..
+  • $S_{\max,r}$ total capacity of ramping up units (ramping down generators will be off before the contingency)
+  • $S_{\min,r}$ minimum working capacity of ramping up units
   • ${D}_{\min,72}$ and ${D}_{\max,72}$: minimum and maximum volumes of predicted demand in next 72 hours.
@@
 2:  Update rolling 24 hour generator states (use them in steps 3-8).
-3:  Increment the number of ramping periods for all ramping generators.
-4:  Switch generators that have finished preparation to ramping up.
-5:  Switch generators that have finished ramping up to on.
-6:  Switch generators that have finished ramping down to off.
-7:  Find must run generators (those whose minimum on time has not been reached or are producing too far above $P_{\min,i}$ to shut off).
-8:  Find can’t start generators (those who have exceeded the maximum allowed daily starts or are already ramping).
-9:  Solve Eq. 6 with $u_{i}\in\{0,1\}$ relaxed to $0\leq y_{i}\leq 1$.
-10:  for Different sets of discretionary generators sufficient to meet the constraints (using Eq. 3 and Eq. 4) do
-11:   Solve economic dispatch Eq. 5 with all must run generators and this discretionary generator set.
-12:   if Objective is lowest so far then
-13:    Update best solution and objective.
-14:   end if
-15:  end for
-16:  return Best solution and best objective.
-17:end for
+3:  Increment ramping periods for ramping units
+4:  Switch unit that finished preparation to ramping up
+5:  Turn units that have finished ramping up to on
+6:  Turn units that have finished ramping down to off
+7:  Find must run units (with minimum on time not yet reached or with $P_{t-1,i}-P_{min,i}$ too large to shut off)
+8:  Find can’t start generators (those that have exceeded maximum allowed daily starts or are already ramping)
+9:  Solve Eq. 6 with $u_{i},v_{i}\in\{0,1\}\rightarrow y_{i},z_{i}\in[0,1]$
+10:  Order units in $\mathcal{G}_{d}$ into sets descending by $y_{i}$, then $z_{i}$
+11:  for Each $\mathcal{G}_{d}$ satisfying (Eq. 6b)-(Eq. 6f) do
+12:   Solve economic dispatch Eq. 5 with $\mathcal{G}_{m},\mathcal{G}_{d}$
+13:   if Objective is lowest so far then
+14:     Update best solution and objective
+15:   end if
+16:  end for
+17:  save Best solution and best objective
+18:end for
 IV-B Simulation and Results
@@
 While ramp characteristics usually come in MW/min, we use the % of $P_{\max,i}$ per minute as the original input into the simulation.
-The percentage values are compiled from several sources, including the academic reports from Deutsches Institut für Wirtschaftsforschung [3] and manufacturer catalog of GE Vernova [13].
+We compile the percentage values the from the academic reports of Deutsches Institut für Wirtschaftsforschung [3] and the manufacturer catalog of GE Vernova [13].
 The provided numbers, although summarising different sources, may still present the upper bound of ramp rates, as comprehensive survey of large thermal units suggest that in real systems power plants do not always deliver the nominal ramp rates specified for equipment [25].
@@
 Once the ramp percentages were assigned to specific generating units, the $P_{\max,i}$ of those generators was multiplied by the ramp rate to get the MW/min value, which was later used in the simulation.
+If the system starts with all generators off, Eq.
+6 is usually infeasible.
+We initialized the system with the fastest ramping generators that can supply the previous demand.
+This may be suboptimal, but it is not the focus of this paper.
+Operators can initialize the system based on their own cold start protocol to a state that is known to be stable.
+If operators are warm starting using Algorithm 2 they can use the current system state.
 TABLE II: Ramp Assumptions and Data Sources
@@
 These results are comparable to those found in Section III using the same software and hardware.
-The runtime scaled roughly as $n^{1.6}$, where $n$ is the number of generators.
-Ideal scaling is believed to be bounded below by $n^{1.5}$, due to the scaling of the underlying linear system solvers [20].
+The runtime scaled roughly as $O(n^{1.6})$, where $n$ is the number of generators.
+Ideal scaling is believed to be bounded below by $O(n^{1.5})$, due to the scaling of the underlying linear system solvers [20].
 This suggests that our optimization solver has close to ideal scaling.
@@
 They are both part of a deterministic shift from the generator being off to on.
-Figure 5: A typical production cycle for generator $i$ in Section V.
-When the generator is on, it can change its production power between $P_{\min,i}$ and $P_{\max,i}$, provided it does not exceed its ramping power.
-It can also turn off if it is close enough to $P_{\min,i}$ compared to its ability to ramp down.
-The only other decision can be made when the generator is off, start ramping up.
-During all other periods, the generator is uncontrollable.
+Figure 5: Unit $i$’s typical production cycle in Section V.
+When on, $P_{i}\in[P_{\min,i},P_{\max,i}]$, though it cannot exceed its ramping power.
+It can turn off if $P_{t-1,i}-r_{d,i}\leq P_{\min,i}$.
+The only other decision can be made when the unit is off, to start ramping up.
+During other periods, it is uncontrollable.
 We introduce a model extension in Fig.
@@
 VI Conclusion and Further Work
-We proposed an algorithm to solve unit commitment with mechanical and ramping constraints by relaxing the integer constraints and then rounding solutions to satisfy the constraints.
-Our algorithm shows between $\approx O(n^{1.6})$ time scaling in the number of generators.
+We proposed an algorithm to solve unit commitment with mechanical and ramping constraints by relaxing the integer constraints and then rounding solutions to satisfy all constraints.
+Our algorithm shows $\approx O(n^{1.6})$ time scaling in the number of generators.
 As the problem size increases, it retains accuracy nearly perfectly.